on a plane, two men together had 135 kilograms of luggage.The first paid $1.35 for his excess luggage and the second paid S2.70 for his excess luggage. had all the luggage belonged to one person, the excess luggage charge would have been $8.10. At most how many kilograms of luggage is each person permitted to bring on the plane free of additional charge.
Answer by mananth(16949)

let x kg be the free allowance
for both of them the free allowance = 2x
135-2x is the excess luggage
135-2x they paid $4.05
1 kg they pay how much ?
They pay 4.05/(135-2x) per kg excess
Single person
135-x is the excess luggage
For 135 kg he pays $8.10
He pays 8.10/(135-x)per kg for excess luggage
The rate per kg for excess luggage is the same so equate both
4.05/(135-2x)=8.10/(135-x)
4.05(135-x)=8.10(135-2x)
546.75-4.05x=1093.5-16.2x
12.15x=546.75
/12.15
x=45 kg free luggage allowance
m.ananth@hotmail.ca